NEW LASER LINES AND FREQUENCY MEASUREMENTS OF FULLY DEUTERATED ISOTOPOMERS OF METHANOL

\underline{E.C.C. VASCONCELLOS}, S.C. ZERBETTO, Instituto de Fisica Gleb Wataghin, Departamento de Eletronica Quantica, Universidade Estadual de Campinas (UNICAMP), 13083-970, Campinas, SP, Brazil; L.R. ZINK, and K.M. EVENSON, National Institute of Standards and Technology, Time and Frequency Division, Boulder, CO 80303-3328.

We report new far-infrared laser lines generated from two fully deuterated isotopomers of methanol: ^{12}CD_{3}OD and ^{13}CD_{3}OD; five from ^{12}CD_{3}OD and 22 from ^{13}CD_{3}OD, which added to the previously reported far-infrared laser lines in these two molecules sums up to 300. The laser lines were mainly pumped by the 10R (regular) and 10SR (sequence) bands of a cw - CO_{2} laser. We measured their frequencies, relative intensities, relative polarizations and pump offsets. The laser lines are in the spectral range from 43.697 to 719.426 μm (6 860 664.6 to 416 710.4 MHz).